# Not All Metrics Are Guilty: Improving NLG Evaluation by Diversifying References

### v2

@@ 3.1 NLG Evaluation Formulation
where $\hat{\mathbf{y}_{0}}=\mathbf{y}^{*}$ and $\mathcal{F}$ is a function leveraged to aggregate scores of multiple diversified sequences, which can be the operation of max aggregation or mean aggregation.
@@ 3.2.1 Paraphrasing Prompt
Following existing work Bawden et al. (2020b), we provide the LLM with the paraphrasing prompt “Paraphrase the sentences: {reference}” to wrap the given reference and employ nucleus sampling Holtzman et al. (2020) to generate a variety of rephrased sentences. In our preliminary experiments, we apply the paraphrasing prompt to paraphrase ten sentences for each English reference sentence from the WMT22 Metrics Shared Task Freitag et al. (2022). We calculate a semantic diversity score [^1] We calculate the mean cosine distance between each rephrased pair using OpenAI Embeddings text-embedding-ada-002. Then, we average the score of each instance to obtain an overall semantic diversity score. of the rephrased sentences as 0.032. We further observe that rephrased sentences primarily involve word-level substitutions, with minimal modifications to the sentence structure. d
@@ 4.1.1 Benchmarks
We choose three meta evaluation benchmarks covering multilingual and multimodal scenarios. These metric benchmarks consist of human scores of the generated text (i.e., $\mathcal{H}(\mathbf{y}^{\prime}|\mathbf{x},\mathbb{Y})$), and we can calculate their correlation with the automatic metric scores $\mathcal{M}(\mathbf{y}^{\prime}|\mathbf{x},A(\mathbb{Y}))$ using multiple references.
@@ 4.1.2 Metrics
We evaluate a variety of automatic metrics covering different categories. Based on the taxonomy of existing work Sai et al. (2022), we select 17 metrics subdivided into five classes:
@@ 4.1.3 Implementation Details
As for our approach, we utilize the gpt-3.5-turbo-instruct model as the LLM along with the instructions outlined in Section 3.2 to diversify the reference sentences into different expressions. When utilizing the OpenAI API, we set the temperature to 1 and the top_p to 0.9. In Equation 3, we employ the max aggregation and generate 10 diversified sentences (i.e., one for each instruction). We further analyze these hyper-parameters in Section 4.3.
@@ 4.2.1 Evaluation on Machine Translation
As shown in the figure 1, our Div-Ref method has shown consistent correlation improvements across all evaluation on the system-level accuracy when compared to the single-reference of the baseline system. Surprisingly, the SOTA metric GEMBA can still be enhanced when evaluated with more references. In terms of different languages, we observe that the diversifying methods are effective across different languages. English and Russian references benefit more than the German ones, which may be due to the distinct multilingual ability of gpt-3.5-turbo. Notably, our approach showcases significant effects on the traditional BLEU metric, which can further facilitate the application due to its efficiency and universality. The large improvement further demonstrates the automatic metric may be not guilty but the evaluation benchmark needs more references.
@@ 4.3 Ablation Analysis
(3) Thirdly, we investigate the aggregation functions using the mean aggregation and the built-in multi-reference aggregation of BLEU and ChrF. We discover that when changing the aggregation from max to mean, the correlation scores for most metrics have dropped, especially in the Chinese-to-English direction. This indicates that the highest-quality reference plays a dominant role in generation evaluation, and our approach to increasing the number of references significantly strengthens this probability. However, averaging multiple reference scores could introduce noise from low-quality reference scores. As for the built-in method of BLEU and ChrF, their performances are indistinguishable.

### v3

@@ 3.1 NLG Evaluation Formulation
where $\hat{\mathbf{y}_{0}}=\mathbf{y}^{*}$ and $\mathcal{F}$ is a function leveraged to aggregate scores of multiple diversified sequences, which can be the operation of maximum aggregation or mean aggregation.
@@ 3.2.1 Paraphrasing Prompt
Following existing work Bawden et al. (2020b), we provide the LLM with the paraphrasing prompt “Paraphrase the sentences: {reference}” to wrap the given reference and employ nucleus sampling Holtzman et al. (2020) to generate a variety of rephrased sentences. In our preliminary experiments, we apply the paraphrasing prompt to paraphrase ten sentences for each English reference sentence from the WMT22 Metrics Shared Task Freitag et al. (2022). We calculate a semantic diversity score [^1] We calculate the mean cosine distance between each rephrased pair using OpenAI Embeddings text-embedding-ada-002. Then, we average the score of each instance to obtain an overall semantic diversity score. of the rephrased sentences as 0.032. We further observe that rephrased sentences primarily involve word-level substitutions, with minimal modifications to the sentence structure.
@@ 4.1.1 Benchmarks
We choose three meta evaluation benchmarks covering multilingual and multimodal scenarios. These metric benchmarks consist of human scores of the generated text (i.e., $\mathcal{H}(\hat{\mathbf{y}}|\mathbf{x},\mathbb{Y})$), and we can calculate their correlation with the automatic metric scores $\mathcal{M}(\hat{\mathbf{y}}|\mathbf{x},A(\mathbb{Y}))$ using multiple references.
@@ 4.1.2 Metrics
We evaluate a variety of automatic metrics covering different categories. Based on the taxonomy of existing work Sai et al. (2022), we select 16 metrics subdivided into five classes:
@@ 4.1.3 Implementation Details
As for our approach, we utilize the gpt-3.5-turbo-instruct model as the LLM along with the instructions outlined in Section 3.2 to diversify the reference sentences into different expressions. When utilizing the OpenAI API, we set the temperature to 1 and the top_p to 0.9. In Equation 3, we employ the maximum aggregation and generate 10 diversified sentences (i.e., one for each instruction). We further analyze these hyper-parameters in Section 4.3.
@@ 4.2.1 Evaluation on Machine Translation
As shown in the Figure 1, our Div-Ref method has shown consistent correlation improvements across all evaluation when compared to the single-reference baseline. Surprisingly, the SOTA metric GEMBA can still be enhanced when evaluated with more references. In terms of different languages, we observe that the diversifying methods are effective across different languages. English and Russian references benefit more than the German ones, which may be due to the distinct multilingual ability of gpt-3.5-turbo-instruct. Notably, our approach showcases significant effects on the traditional BLEU metric, which can further facilitate the application due to its efficiency and universality. The large improvement further demonstrates the automatic metric may be not guilty but the evaluation benchmark needs more references.
@@ 4.3 Ablation Analysis
(3) Thirdly, we investigate the aggregation functions using the mean aggregation and the built-in multi-reference aggregation of BLEU and ChrF. We discover that when changing the aggregation from maximum to mean, the correlation scores for most metrics have dropped, especially in the Chinese-to-English direction. This indicates that the highest-quality reference plays a dominant role in generation evaluation, and our approach to increasing the number of references significantly strengthens this probability. However, averaging multiple reference scores could introduce noise from low-quality reference scores. As for the built-in method of BLEU and ChrF, their performances are indistinguishable.